OpenStudy (anonymous):

The horizontal shadow of a vertical tree is 23.4 m long when the angle of elevation of the sun is 56.3 degrees. How tall is the tree? Round to the nearest meter.

OpenStudy (amistre64):

id try to use a tangent function ....

OpenStudy (amistre64):



OpenStudy (anonymous):

How would I do that? :P

OpenStudy (amistre64):

you need to recall the definition for the tangent function, the rest is just basic algebra

OpenStudy (amistre64):

once you have that definition, let me know

OpenStudy (anonymous):

Opposite/Adjacent?

OpenStudy (amistre64):

yep so they give us an angle, a, and a known length, k. so, tan(a) = h/k h = k tan(a) which a calculator set to degree mode should pop out

OpenStudy (anonymous):

So h=23.4*tan(56.3) h=35.0868 Right?

OpenStudy (amistre64):

thats what i get ... so rnding to 35m should suffice

OpenStudy (anonymous):

Ok :) thank you.

OpenStudy (amistre64):

youre welcome